OpenStudy (anonymous):

Is (y+1)^2 - 3 = 2x a function? If it is, what is its domain and range?

OpenStudy (anonymous):

ya it is a function it fefines implicitly f(x) = y in some domain (have to ckeck carefully) becouse F(x,y) has not vanishing and continuous partial derivative of y.

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

what u study in, university, highschool?

OpenStudy (anonymous):

high school

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

then u can wtite it like this: \[((y+1)^{2}-3)/2=x\] then it becomes function of y, it means x =f(y) it is well defind function, (inverse of y) with domain all R, and range (-3/2,infinity)

OpenStudy (anonymous):

for finding f(x) = y, make change of variable y' = y+1 so your exprecion becomes\[y'^{2} = 2x+3\] which is equal to\[y' =\sqrt{2x +3}\] domain of this is x>-3/2 and range all positive real numbers

OpenStudy (anonymous):

ah, forgot one thing: remmeber to come back to your variable y = y' -1, so range will be -1,infinity.

OpenStudy (anonymous):

hope it helps

OpenStudy (anonymous):

so it goes: \[(y+1)^{2} - 3 = 2x\] \[(y+1)^{2}-3-2x = 0\] \[(y+1)^{2} = 3+2x\] \[y = -1 +/- \sqrt{3+2x}\]

OpenStudy (anonymous):

but what happens again for x?

OpenStudy (anonymous):

the exprecion under the root sign can't be negative, so x cant be smaller than -3/2

OpenStudy (anonymous):

it means the domain is (-3/2,infinity)